OpenStudy (anonymous):

Solve for x: 5 − (x + 5) > −2(x + 4) x > −8 x < −8 x > −18 x < −18

Nnesha (nnesha):

distribute left bracket by -1 and right side bracket by -2 do you know how do distribute ???

Nnesha (nnesha):

multiply everything which is in bracket by outside number for example a(b-c) = ab -ac

OpenStudy (anonymous):

oh okay so 5*5(left) and and 2*4 (right)?

Nnesha (nnesha):

not 5 you just have to multiply by -1 \[- (x +5) = -1(x) , -1(5) = ??\]

OpenStudy (anonymous):

-5

Nnesha (nnesha):

-1 times x = ??

OpenStudy (anonymous):

-1x

Nnesha (nnesha):

yes right now distribute right side by -2

Nnesha (nnesha):